OpenStudy (anonymous):

Please help: (Proving the Euclidean algorithm maybe?) Here's the question: Show that if a

OpenStudy (kinggeorge):

Start with \(g=\gcd(a, b)\). This means that \(g|a\) and \(g|b\). By some properties of divisible, you have that \(g|k\,a\) where \(k\) is an integer. Also, this means that \(g|b-k\,a\) since if a number divides two other numbers, it divides their difference.

OpenStudy (kinggeorge):

Thus, \(g|a\) and \(g|(b-k\,a)\). Since \(0<b-k\,a<b\) this means that \(g\) is also the gcd of a and b-ka.

OpenStudy (kinggeorge):

I suppose I should have put \(0\leq b-k\,a<a\) instead.

OpenStudy (anonymous):

So would this be correct: Show that if a<b, then gcd(a,b) = gcd(a, b-ka), where 0≤ b-ka < a Let g = gcd(a,b) (Let m = gcd(a, b-ka) Show g = m) DO I NEED THIS PART??? We know that g|a and g|b So g| (b-a) We know that g|ka for some integer k, and g|b So g| (b-ka) OR USING M... We know that m|a and m|b-ka So m| (a + (b - ka)) So m| a (b-k) See I'm lost on how to say it... :( Therefore g|a and g| (b-ka)

OpenStudy (kinggeorge):

The way you're trying to do it would work if you were to show that \(m\leq g\) and that \(g\leq m\).

OpenStudy (anonymous):

Here's what our teacher gave us to start from and told us to modify it for this new problem. He said: If a <b, then gcd(a, b) = gcd (a, b-a) Let d= gcd(a,b) Let m = gcd (a, b-a) Show d= m We know that d|a and d|b So d| (b-a) Therefore d|a and d| (b-a) So d ≤ m We know that m|a and m| (b-a) So m| (a + b - a) d| b Therefore m|a and m|b So m ≤ d So m = d I need to modify all of that for "If a <b, then gcd (a, b) = gcd (a, b-ka), where 0 ≤ b - ka < a "

OpenStudy (kinggeorge):

So you would have something like this. Let g= gcd(a,b) Let m = gcd (a, b-ka) Show g= m We know that g|a and g|b so g|ka, So g| (b-ka) Therefore g|a and g| (b-ka) So g ≤ m We know that m|a and m| (b-ka) So m| (a + b - ka) so m|(a(1-k)+b) Since m|a, m|(a(1-k)) Thus, m| b Therefore m|a and m|b So m ≤ g So m = d

OpenStudy (kinggeorge):

There's just a few extra steps in there to show that \(m\leq g\)

OpenStudy (anonymous):

you put in the extra steps? or should I figure those out...

OpenStudy (kinggeorge):

I already put the extra steps in the explanation.

OpenStudy (kinggeorge):

Although it should say \(m=g\) at the end. Not m=d

OpenStudy (anonymous):

ok thanks, so should that last part say We know that m|a and m| (b-ka) So m| (a + b - ka) so m|(a(1-k)+b) Since m|a, m|(a(1-k)), and m| b Therefore m|a and m|b So m ≤ g So m = d Guess I'm confused how you can say Thus m| b

OpenStudy (kinggeorge):

Suppose \(m \nmid b\). Since \(m|a(1-k)\) we can write it as \(a(1-k))=mz\) where \(z\in\mathbb{Z}\). This means that \(m|mz+b\) so we need to be able to factor an \(m\) our of \(mz+b\). If \(m \nmid b\) then it's impossible! This is a contradiction, so \(m|b\).

OpenStudy (anonymous):

Ok, so I should include that contradiction in the problem, or is it generally understood?

OpenStudy (kinggeorge):

That's generally understood. I probably wouldn't include it in one of my proofs.

OpenStudy (anonymous):

Gotcha. Thank you so much for your help, you are awesome :D

OpenStudy (kinggeorge):

You're welcome.